. What is the equation for Boylès LaW Solve the following problems: At 103 torr a gas has a volume of 5.2 L. What is the volume if the pressure is increased to 400 torr? a. b. A gas at 0.85 atm has a volume of 200 ml.. What pressure will make the volume 950 mL? of an ideal gas exnlain why pressure
Answer #1

4.

By Boyle's law we know that at constant temperature and equal moles the product of pressure and volume remains the same. So,

PV = constant, so, P1V1 = P2V2

Here P represents the pressure and V is the volume of the gas.

5.

Applying this equation P1V1 = P2V2 in the part a we have

103 torr * 5.2 L = 400 torr * V

where V reprsents the volume when pressure is 400 torr

So, V = 103 * 5.2 / 400 = 1.339 L

For part b we have

0.85 atm * 200 ml = P * 950 ml

so, P = 0.85*200/950 atm = 0.179 atm